OpenStudy (anonymous):

a glass prism used in the study of optics has a right triangular base. Th legs of the triangle are 3.00cm and 4.00cm. The prism is 8.50cm high. what is the total surface area of the prism?

OpenStudy (anonymous):

i did 4*8.5+3*8.5+12*8.5=102 is this correct? and what to do next?

OpenStudy (anonymous):

The prism has 5 sides, we must find the area of each one. 2 of the sides will be called the "ends" that you can see in the image. They are calculated simply by the equation of a triangle: A=(1/2)*b*h A=(1/2)*4*3=6 cm squared We multiply this value by 2 because there are two "ends" to the prism so Aends= 2*6= 12 cm squared Now we must find the area of the rectangular pieces of the prism: A1=3*8.5=22.5 cm squared A2=4*8.5=34 cm sqared The last side must use the Pythagorean theorem to find the length: \[\sqrt{3^{2}+4^{2}}=5\] A3=8.5*5= 42.5 cm squared Adding all the areas gives you: Aends +A1 +A2 +A3= 111 cm squared

OpenStudy (anonymous):

thanks for explaining